# spacing=nonfrench

### v1

@@ Abstract
We conduct extensive experiments on two autonomous driving benchmark datasets to evaluate the robustness of our model against non-causal perturbations and its generalization capacity. Our results indicate that the robustness of predictions can be improved by up to $\mathbf{54\%}$ without a significant detriment to prediction accuracy. Lastly, we demonstrate the superior domain generalizability of the proposed model, which achieves up to $\mathbf{29\%}$ improvement in cross-domain performance. These results underscore the potential of our model to enhance both robustness and generalization capacity for trajectory prediction in diverse autonomous driving domains. [^2] The project page is available at: http://critic-model.github.io
@@ I Introduction
MTR [2]
Refer to caption Refer to caption
CRiTIC (Ours)
Refer to caption Refer to caption
@@
Figure 2: An overview of CRiTIC. In this architecture, Causal Discovery Network receives the map-aware agent representations and generates a causality adjacency matrix. The matrix is used by a Transformer-based prediction backbone to shape the attention toward the causal agents.
@@ III Problem Formulation
Trajectory Prediction. The goal is to predict future trajectories (in $XYZ$ coordinates) of target agents $\mathbf{T}\in\mathbb{R}^{N_{\text{targ}}\times M\times T_{f}\times 3}$ (where $M=6$ and $T_{f}$ denote the number of modes and future time steps, respectively) given the observed agents’ trajectories $\mathbf{S}_{A}$ and scene context $\mathbf{S}_{M}$ as: $\mathbf{T}=\textup{F}(\mathbf{S}_{A},\mathbf{S}_{M})$. The observed trajectories $\mathbf{S}_{A}\in\mathbb{R}^{N_{\text{obs}}\times T_{h}\times D_{a}^{0}}$ (where $T_{h}$ is the number of history time steps) include features, such as coordinates, heading angle, velocity, dimensions, and agent type. The scene context $\mathbf{S}_{M}\in\mathbb{R}^{N_{m}\times N_{p}\times D_{m}^{0}}$ represents map components (e.g., lanes, road boundaries) as $N_{m}$ polylines characterized by position, component type, and dynamic states.
@@
For causal trajectory prediction, we adopt the formulation of the amortized causal discovery (ACD) approach, which builds upon the principles of Granger causality [17]. Its key assumption is that for a set of data samples $\mathbf{x}_{s}$, with a varying set of summary causal graphs, $\mathcal{G}_{s}$, there exists a shared dynamics model $g$ that $\mathbf{x}^{t+1}_{s}=g(\mathbf{x}_{s}^{\leq t},\mathcal{G}_{s})+\bm{\varepsilon}_{s}^{t+1}$. Here, $\mathcal{G}_{s}$ represents the causal graph for sample $s$, $\mathbf{x}_{s}\in\mathbb{R}^{N_{\text{s}}\times T}$ denotes $N_{s}$ time series of length $T$, and $\varepsilon_{s}^{t+1}$ is the independent noise term. This formulation allows us to separate the data generation process into two components and express the process as: $\mathbf{x}^{t+1}_{s}\approx f_{\theta}(\mathbf{x}_{s}^{\leq t},f_{\phi}(\mathbf{x}_{s}^{\leq t}))$, with the components: (i) causal discovery model $f_{\phi}$ —this is the CDN module explained in Section IV-B— which infers the causal structure from the observed data, and (ii) shared dynamics model $f_{\theta}$ —this is the trajectory prediction backbone, $\mathbf{T}$, discussed in Section IV-C— which predicts future states based on past observations and the inferred causal structure.
@@ IV-A AgentNet: Generating Agent-Centric Representations
AgentNet is composed of two submodules: trajectory encoder and map-encoder. The context information generated by the map encoder is fused with the outputs of the trajectory encoder to form the contextual map representations. Note that in this stage the agents’ states are encoded independently, as we need disentangled agent representations for explicit causal interaction modeling via the CDN.
Map Encoder. First, we encode each map component using a PointNet-based encoder [47], followed by a max-pooling layer, $\mathbf{S}^{\prime}_{M}=\textup{MaxPool(PFFN}(\mathbf{S}_{M}))$, where PFFN is a point-wise feed-forward network, $\mathbf{S}^{\prime}_{M}\in\mathbb{R}^{N_{m}\times D_{m}^{1}}$ is the matrix of map features, and $D_{m}^{1}$ is the dimensions of the map features. To efficiently incorporate map information into agent representations, we employ a multi-context gating (MCG) layer [1] to gather information from $N_{m}$ map elements into a concise map context vector $\mathbf{S}^{\prime\prime}_{M}\in\mathbb{R}^{D_{m}^{2}}$, where $\mathbf{S}^{\prime\prime}_{M}=\textup{MCG}(\mathbf{S}^{\prime}_{M})$.
Context-aware Trajectory Encoder. The agents’ states, while presented in the target agent’s coordinate frame, are processed using a Pointnet-based encoder [47], followed by a max-pooling and a GRU layer [48] used to encode the Kinematics of the agents. Ultimately, integrating the map context vector into the agent representation ensures the inclusion of map features, resulting in what we refer to as map-aware agent representations. To get intermediate temporal trajectory representations we have $\mathbf{S}^{\prime}_{A}=\textup{MLP}(\mathbf{S}_{A})$, where $\mathbf{S}^{\prime}_{A}\in\mathbb{R}^{N_{obs}\times T\times D_{a}^{1}}$. Then, $\mathbf{S}^{\prime\prime}_{A}=\textup{GRU}(\textup{MaxPool}(\mathbf{S^{\prime}}_{A}))\oplus\mathbf{S}^{\prime\prime}_{M}$, where $\oplus$ represents the concatenation operator and $\mathbf{S}^{\prime\prime}_{A}\in\mathbb{R}^{N_{obs}\times D_{a}^{2}}$ is the matrix of the agent representations.
@@ IV-B Causal Discovery Network (CDN)
We capture the inter-agent causal relations in the form of a directed graph, where the vertices are the agent representations and the edges indicate the causal relations among the agents. The role of CDN is to process the set of map-aware agent representations generated by AgentNet and identify the structure of the causal summary graph. In other words, CDN determines whether for each pair of agents $(i,j)$, whether agent $i$ has a causal influence on agent $j$.
@@
$\begin{split}\mathbf{s}^{1}_{i}&=\gamma^{0}(\mathbf{s}^{0}_{i}),\\ \mathbf{s}^{2}_{i}&=\gamma^{1}\left(\mathbf{s}^{1}_{i},\sum_{j\in\mathcal{N}(i)}\phi^{1}\left(\mathbf{s}^{1}_{i},\mathbf{s}^{1}_{j}\right)\right),\\ e_{ij}&=\left\{\begin{matrix}\rho(\mathbf{W_{\text{MPNN}}}^{T}\phi^{2}(\mathbf{s}_{i},\mathbf{s}_{j}),\lambda)&i\neq j\\ 1&i=j\end{matrix}\right.,\\ \end{split}$ (1)
where $\mathbf{s}^{0}_{i}=\mathbf{S}^{\prime\prime}_{M}$ is the initial representation of the agent $i$ generated by AgentNet, $\mathcal{N}(i)$ is the set of neighbors of node $i$ (assumed to have an initial fully-connected graph without self-loops), $\gamma^{0},\gamma^{1},\phi^{1},\phi^{2}$ are two-layer MLP networks, and $\mathbf{W}_{\text{MPNN}}$ is a learnable linear projection weight.
The desired set of edges is discrete random variables from the Bernoulli distribution. However, training models with discrete variables is challenging due to the non-differentiability of the sampling operation. To overcome this issue, we replace the binary edge variables with a low-variance continuous relaxation of it named “BinConcrete” [50], which is represented by the function $\rho$ in the above equation. It is defined as: $\text{BinConcrete}(\alpha,\lambda)\doteq\text{Sigmoid}((L+\log\alpha)/\lambda)$, where $\lambda$ is the temperature hyperparameter and $L=\log U+\log(1-U)$, where $U$ is a random variable sampled from a uniform distribution. Finally, the weighted edges $e_{ij}\in[0,1]$ form the weighted adjacency matrix $\mathbf{A}$.
@@
Sparsity Regularization. A possible degenerate state for the causal discovery network is always generating a densely connected graph, or in the extreme case, a fully connected graph. Obviously, such a graph is not able to discriminate causal vs. non-causal relations. Inspired by the sparsity of causal connections among the agents [5], to discourage this phenomenon, we add an adjacency matrix sparsity regularization loss by defining it as the KL divergence between the marginal probability of an edge being causal and a fixed prior binary distribution with the hyperparameter $p$ indicating the marginal probability for having a causal link.
@@
Auxiliary Denoising Autoencoder (DAE). Following the definition of the Granger causality for time series data in Section IV, the causal graph aids the prediction of future variables from the past value of its parents. Motivated by this we add the DAE task as an auxiliary supervision to facilitate the causal discovery. In this task, the objective is to reconstruct the values of the masked intermediate temporal agent representations generated by AgentNet based on the values of the other vertices and the causal graph. Note that, using temporal features for this task best matches the definition of Granger causality.
@@
Thereby, we employ a two-layer graph convolutional network (GCN) as a denoising autoencoder (DAE), where the graph is defined as: $\mathcal{G}_{\text{DAE}}=\{\mathcal{V}_{\text{DAE}},\mathcal{E}_{\text{DAE}}\}$, the vertices are $\mathcal{V}_{\text{DAE}}=\{v_{i}^{t}|v_{i}^{t}\in\mathbf{H}^{0},i\in\{1,\dots,N\},t\in\{1,\dots,T^{\prime}\}\}$, where for computational efficiency we downsample and reshape the temporal agent representations, $\mathbf{H}^{0}=\text{Reshape}(\text{AvgPool}(\text{SG}(\mathbf{S}^{\prime}_{A})))\in\mathbb{R}^{N_{DAE}\times D_{at}}$, where $N_{\text{DAE}}=T^{\prime}\times N_{\text{obs}}$, and $T^{\prime}=3$ is the downsampled temporal dimension. To avoid the model collapse to naïve solutions, we detach the gradients using the Stop Gradient operation denoted by SG so the DAE loss cannot directly affect representation learning in AgentNet. The edges are defined as $\mathcal{E}_{\text{DAE}}=\{(v_{i}^{t},v_{j}^{t^{\prime}})|t\leq t^{\prime},(e_{ij}=1\;\text{or}\;i=j)\}$. The edges $\mathcal{E}_{\text{DAE}}$ correspond to the adjacency matrix $\mathbf{A}_{\text{DAE}}$, which is a block lower-triangular extension of the adjacency matrix generated by the CDN.
@@
Next, we mask a random selection of vertices using a binary mask $\mathbf{M}\in\mathbb{R}^{N_{\text{DAE}}\times D_{at}}$ controlled by the masking ratio $r$. The masked representation is given by $\mathbf{H}^{1}=\mathbf{H}^{0}\odot\mathbf{M}$, where $\odot$ is the Hadamard product operator. We constrain the mask to have an all-equal last dimension, i.e., we perform vertex-wise masking. Subsequently, the GCN layers are defined as: $\mathbf{H}^{2}=\textup{ReLU}\left(\mathbf{B}\mathbf{H^{1}}\mathbf{W^{1}}\right),\;\mathbf{H}^{3}=\mathbf{A}_{\text{DAE}}\mathbf{H}^{2}\mathbf{W^{2}},$ where $\mathbf{W}^{1}$ and $\mathbf{W}^{2}$ are the linear projection weight matrices, and $\mathbf{H}^{2}$ and $\mathbf{H}^{3}$ denote the intermediate vertex representations, and the reconstructed agent representations, respectively. Loss function $L_{\textup{DAE}}$ is equal to negative cosine similarity between masked vertices of $\mathbf{H}^{0}$ and $\mathbf{H}^{3}$.
@@ IV-C Causal Trajectory Prediction
where $\mathbf{\Phi}=\textup{Softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})$, $A^{C}=\mathbf{J}-\mathbf{A}$, $\mathbf{J}\in\mathbb{R}^{N_{obs}\times N_{obs}}$ is an all-ones matrix, $\mathbf{V}^{\prime}$ denotes the column-normalized value matrix in attention layers, $\mathbf{N}=\mathcal{N}(\mathbf{0},\mathbf{1})$ is a Gaussian noise matrix of the shape $\mathbb{R}^{N_{obs}\times d_{v}}$, $d_{v}$ is the dimension of the value vector, and $\alpha$ is the noise scale hyperparameter, which is set to zero at the inference time.
@@
Backbone Network. The backbone network, MTR [2], is composed of 6 layers of Transformer encoder blocks, a dense prediction (DFP) network, and 6 layers of decoder Transformer blocks with latent anchors and dynamic map query layers. Unlike the original backbone model, we only feed the map-aware agent representations to the Encoder network and apply CAG for its self-attention layers (with global attention). We also apply the CAG for the anchor-agent cross-attention layers of the decoder. Due to the limited space, we refer the reader for additional details regarding the backbone to [2].

### v2

@@ Abstract
We conduct extensive experiments on two autonomous driving benchmark datasets to evaluate the robustness of our model against non-causal perturbations and its generalization capacity. Our results indicate that the robustness of predictions can be improved by up to $\mathbf{54\%}$ without a significant detriment to prediction accuracy. Lastly, we demonstrate the superior domain generalizability of the proposed model, which achieves up to $\mathbf{29\%}$ improvement in cross-domain performance. These results underscore the potential of our model to enhance both robustness and generalization capacity for trajectory prediction in diverse autonomous driving domains. [^4] The project page is hosted at: http://ehsan-ami.github.io/critic
@@ I Introduction
CRiTIC (Ours)                        MTR [2]
@@
Figure 2: An overview of CRiTIC. In this architecture, Causal Discovery Network receives the agent representations and generates a causality adjacency matrix. The matrix is used by a Transformer-based prediction backbone to shape the attention toward the causal agents.
@@ III Problem Formulation
Trajectory Prediction. The goal is to predict future trajectories (in $XYZ$ coordinates) of target agents $\mathbf{T}\in\mathbb{R}^{N_{\text{targ}}\times M\times T_{f}\times 3}$ (where $N_{\text{targ}}$ represents the number of target agents, $M=6$ and $T_{f}$ denote the number of modes and future time steps, respectively) given the observed agents’ trajectories $\mathbf{S}_{A}$ and scene context $\mathbf{S}_{M}$ as: $\mathbf{T}=\textup{F}(\mathbf{S}_{A},\mathbf{S}_{M})$. The observed trajectories $\mathbf{S}_{A}\in\mathbb{R}^{N_{\text{obs}}\times T_{h}\times D_{a}^{0}}$ (where $N_{\text{obs}}$ is the number of observed agents in the scene, $T_{h}$ is the number of history time steps, and $D_{a}^{0}$ is the dimension of agent features) include features, such as coordinates, heading angle, velocity, dimensions, and agent type. The scene context $\mathbf{S}_{M}\in\mathbb{R}^{N_{m}\times N_{p}\times D_{m}^{0}}$ represents map components (e.g., lanes, road boundaries) as $N_{m}$ polylines, each consisting of up to $N_{p}$ points, and characterized by position, component type, and dynamic states with the total feature dimension of $D_{m}^{0}$.
@@
For causal trajectory prediction, we adopt the formulation of the amortized causal discovery (ACD) approach, which builds upon the principles of Granger causality [17]. Its key assumption is that for a set of data samples $\mathbf{x}_{s}$, with a varying set of summary causal graphs, $\mathcal{G}_{s}$, there exists a shared dynamics model $g$ that $\mathbf{x}^{t+1}_{s}=g(\mathbf{x}_{s}^{\leq t},\mathcal{G}_{s})+\bm{\varepsilon}_{s}^{t+1}$. Here, $\mathcal{G}_{s}$ represents the causal graph for sample $s$, $\mathbf{x}_{s}\in\mathbb{R}^{N_{\text{s}}\times T}$ denotes $N_{s}$ time series of length $T$, and $\varepsilon_{s}^{t+1}$ is the independent noise term modeling inherent aleatoric uncertainty of the data. This formulation allows us to separate the data generation process into two components and express the process as: $\mathbf{x}^{t+1}_{s}\approx f_{\theta}(\mathbf{x}_{s}^{\leq t},f_{\phi}(\mathbf{x}_{s}^{\leq t}))$, with the components: (i) causal discovery model $f_{\phi}$ —this is the CDN module explained in Section IV-B— which infers the causal structure from the observed data, and (ii) shared dynamics model $f_{\theta}$ —this is the trajectory prediction backbone, $\mathbf{T}$, discussed in Section IV-C— which predicts future states based on past observations and the inferred causal structure.
@@ IV-A AgentNet: Generating Agent-Centric Representations
AgentNet is composed of two submodules: trajectory encoder and map-encoder. First, the raw map and agent inputs are presented in an agent-centric approach [2, 1]. The context information generated by the map encoder is fused with the outputs of the trajectory encoder to form the contextual agent representations. Note that in this stage the agents’ states are encoded independently, as we need disentangled agent representations for explicit causal agent interaction modeling via the CDN.
Map Encoder. We encode each map component using a PointNet-based encoder [47], followed by a max-pooling layer, $\mathbf{S}^{\prime}_{M}=\textup{MaxPool(PFFN}(\mathbf{S}_{M}))$, where PFFN is a point-wise feed-forward network, $\mathbf{S}^{\prime}_{M}\in\mathbb{R}^{N_{m}\times D_{m}^{1}}$ is the matrix of map features, and $D_{m}^{1}$ is the dimension of the map features. To efficiently incorporate map information into agent representations, we employ a multi-context gating (MCG) layer [1] to gather information from $N_{m}$ map elements into a concise map context vector $\mathbf{S}^{\prime\prime}_{M}\in\mathbb{R}^{D_{m}^{2}}$, where $\mathbf{S}^{\prime\prime}_{M}=\textup{MCG}(\mathbf{S}^{\prime}_{M})$.
Context-aware Trajectory Encoder. The agents’ states, while presented in the target agent’s coordinate frame, are processed using a Pointnet-based encoder [47], followed by a max-pooling and a GRU layer [48] used to encode the Kinematics of the agents. To get intermediate temporal trajectory representations we have $\mathbf{S}^{\prime}_{A}=\textup{MLP}(\mathbf{S}_{A})$, where $\mathbf{S}^{\prime}_{A}\in\mathbb{R}^{N_{obs}\times T\times D_{a}^{1}}$ and $D_{a}^{1}$ is the dimension of the intermediate agent representations. Then, $\mathbf{S}^{\prime\prime}_{A}=\textup{GRU}(\textup{MaxPool}(\mathbf{S^{\prime}}_{A}))\oplus\mathbf{S}^{\prime\prime}_{M}$, where $\oplus$ represents the concatenation operator. Here, the single map context vector $\mathbf{S}^{\prime\prime}_{M}$ is concatenated with each agent’s representation separately. This results in $\mathbf{S}^{\prime\prime}_{A}\in\mathbb{R}^{N_{obs}\times D_{a}^{2}}$, the matrix of map-aware agent representations with dimension $D_{a}^{2}$.
@@ IV-B Causal Discovery Network (CDN)
We capture the inter-agent causal relations in the form of a directed graph, where the nodes are the agent representations and the edges indicate the causal relations among the agents. The role of CDN is to process the set of map-aware agent representations generated by AgentNet and identify the structure of the causal summary graph. In other words, CDN determines whether for each pair of agents $(i,j)$, whether agent $i$ has a causal influence on agent $j$.
@@
$\begin{split}\mathbf{s}^{1}_{i}&=\gamma^{0}(\mathbf{s}^{0}_{i}),\\ \mathbf{s}^{2}_{i}&=\gamma^{1}\left(\mathbf{s}^{1}_{i},\sum_{j\in\mathcal{N}(i)}\phi^{1}\left(\mathbf{s}^{1}_{i},\mathbf{s}^{1}_{j}\right)\right),\\ e_{ij}&=\left\{\begin{matrix}\rho(\mathbf{W_{\text{MPNN}}}^{T}\phi^{2}(\mathbf{s}^{2}_{i},\mathbf{s}^{2}_{j}),\lambda)&i\neq j\\ 1&i=j\end{matrix}\right.,\\ \end{split}$ (1)
where $\mathbf{s}^{0}_{i}\in\mathbf{S}^{\prime\prime}_{A}$ is an individual initial node representation in the MPNN, and it is set to representation vector of the agent $i$ generated by AgentNet. Consequently, $\mathbf{s}^{1}_{i}$ and $\mathbf{s}^{2}_{i}$ are intermediate node representations of agent $i$. The set $\mathcal{N}(i)$ includes neighbors of node $i$ (we assume to have an initial fully-connected adjacency graph without self-loops), $\gamma^{0},\gamma^{1},\phi^{1},\phi^{2}$ are two-layer MLP networks and $\mathbf{W}_{\text{MPNN}}$ is a learnable linear projection weight.
Causal edges could be presented by binary discrete random variables which have Bernoulli distributions. However, training models with discrete variables is challenging due to the non-differentiability of the sampling operation. To overcome this issue, we replace the binary edge variables with a low-variance continuous relaxation of it named “BinConcrete” [50], which is represented by the function $\rho$ in the above equation. It is defined as: $\text{BinConcrete}(\alpha,\lambda)\doteq\text{Sigmoid}((L+\log\alpha)/\lambda)$, where $\lambda$ is the temperature hyperparameter and $L=\log U+\log(1-U)$, where $U$ is a random variable sampled from a uniform distribution. Finally, the weighted edges $e_{ij}\in[0,1]$ form the weighted adjacency matrix $\mathbf{A}$.
@@
Sparsity Regularization. A possible degenerate state for the causal discovery network is always generating a densely connected graph, or in the extreme case, a fully connected graph. Obviously, such a graph is not able to discriminate causal vs. non-causal relations. Inspired by the sparsity of causal interactions among the agents [5], we add an adjacency matrix sparsity regularization loss. This loss is defined as the KL divergence between the marginal probability of an edge being causal and a fixed prior binary distribution with the hyperparameter $p$ indicating the marginal probability for having a causal link. A small value of this parameter creates an information bottleneck by inducing a sparse causal interaction graph. The causally gated backbone prediction model, as explained later in IV-C, has limited information interchange bandwidth among agent representations. Therefore, the model is incentivized to allocate its limited attention capacity to truly causal agents to minimize the loss function.
@@
Auxiliary Denoising Autoencoder (DAE). Following the definition of the Granger causality for time series data in Section IV, the causal graph aids the prediction of future variables from the past value of its parents. Motivated by this we add the DAE task as an auxiliary supervision to facilitate the causal discovery. In this task, the objective is to reconstruct the values of the masked intermediate temporal agent representations generated by AgentNet based on the values of the other nodes and the causal graph. Note that, using temporal features for this task best matches the definition of Granger causality.
@@
Thereby, we employ a two-layer graph convolutional network (GCN) as a denoising autoencoder (DAE), where the graph is defined as: $\mathcal{G}_{\text{DAE}}=\{\mathcal{V}_{\text{DAE}},\mathcal{E}_{\text{DAE}}\}$, the nodes are $\mathcal{V}_{\text{DAE}}=\{v_{i}^{t}|v_{i}^{t}\in\mathbf{H}^{0},i\in\{1,\dots,N\},t\in\{1,\dots,T^{\prime}\}\}$, where for computational efficiency we downsample the temporal agent representations, $\mathbf{H}^{0}=\text{Flatten}(\text{GroupAvgPool}(\text{SG}(\mathbf{S}^{\prime}_{A})))\in\mathbb{R}^{N_{DAE}\times D_{at}}$, where $N_{\text{DAE}}=T^{\prime}\times N_{\text{obs}}$, $T^{\prime}=3$ is the downsampled temporal dimension, and $D_{at}=D_{a}^{1}$ is the feature dimension of the temporal agent representations. Downsampling is done via group average pooling, with groups defined by chunking the temporal dimension. The tensor Flatten operation is then applied so that each node $v_{i}^{t}$ corresponds to a specific agent at a particular chunked temporal state. To avoid the model collapse to naïve solutions, we detach the gradients using the Stop Gradient operation denoted by SG so the DAE loss cannot directly affect representation learning in AgentNet. The edges are defined as $\mathcal{E}_{\text{DAE}}=\{(v_{i}^{t},v_{j}^{t^{\prime}})|t\leq t^{\prime},(e_{ij}=1\;\text{or}\;i=j)\}$. The edges $\mathcal{E}_{\text{DAE}}$ correspond to the adjacency matrix $\mathbf{A}_{\text{DAE}}$, which is a block lower-triangular extension of the adjacency matrix generated by the CDN.
@@
Next, we mask a random selection of nodes using a binary mask $\mathbf{M}\in\mathbb{R}^{N_{\text{DAE}}\times D_{at}}$ controlled by the masking ratio $r$. The masked representation is given by $\mathbf{H}^{1}=\mathbf{H}^{0}\odot\mathbf{M}$, where $\odot$ is the Hadamard product operator. We constrain the mask to have an all-equal last dimension, i.e., we perform node-wise masking. Subsequently, the GCN layers are defined as: $\mathbf{H}^{2}=\textup{ReLU}\left(\tilde{\mathbf{A}}_{\text{DAE}}\mathbf{H^{1}}\mathbf{W^{1}}\right),\;\mathbf{H}^{3}=\tilde{\mathbf{A}}_{\text{DAE}}\mathbf{H}^{2}\mathbf{W^{2}},$ where $\tilde{\mathbf{A}}_{\text{DAE}}\in\mathbb{R}^{N_{\text{DAE}}\times N_{\text{DAE}}}$ is the row-normalized adjacency matrix, $\mathbf{W}^{1}$ and $\mathbf{W}^{2}$ are the linear projection weight matrices, and $\mathbf{H}^{2}$ denotes the intermediate node representations, and $\mathbf{H}^{3}$ is the reconstructed node representations. Loss function $L_{\textup{DAE}}$ is equal to negative cosine similarity between masked nodes of $\mathbf{H}^{0}$ and $\mathbf{H}^{3}$.
@@ IV-C Causal Trajectory Prediction
where $\mathbf{\Phi}=\textup{Softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})$, $A^{C}=\mathbf{J}-\mathbf{A}$, $\mathbf{J}\in\mathbb{R}^{N_{obs}\times N_{obs}}$ is an all-ones matrix, $\mathbf{V}^{\prime}$ denotes the column-normalized value matrix in attention layers, $\mathbf{N}=\mathcal{N}(\mathbf{0},\mathbf{1})$ is a Gaussian noise matrix of the shape $\mathbb{R}^{N_{obs}\times d_{v}}$, $d_{v}$ and $d_{k}$ are the dimensions of the key and value vectors, respectively. The noise scale hyperparameter $\alpha$ is set to zero at the inference time.
@@
Backbone Network. The backbone network, MTR [2], is composed of 6 layers of Transformer encoder blocks, a dense prediction (DFP) network, and 6 layers of decoder Transformer blocks with latent anchors and dynamic map query layers following [2]. In this study we focus on the inter-agent causal discovery, and so unlike the original backbone model, we only feed the map-aware agent representations to the Encoder network and apply CAG for its self-attention layers (with global attention). This modification is only made to isolate and focus on inter-agent causal relationships without interference from map-based interactions. We leave the investigation of agent-map causal interactions as an important direction for future work.